Find an expression of z^n which nEN


1
Expert's answer
2021-07-27T06:12:02-0400

"z =\frac{z_1}{z_2} \\
z_1 = tan θ + i \\
z_2 = \bar{z_1} = \bar{(\tan\theta+i)}=tan \theta -i\\
z =\frac{z_1}{z_2} =\frac{tan \theta +i}{tan \theta -i}\\
z =\frac{tan \theta +i}{tan \theta -i}*\frac{tan \theta +i}{tan \theta +i}\\
Simplifying \\
z= \frac{tan^2 \theta +i^2+2 tan \theta i}{tan^2 \theta +1}\\
z= \frac{tan^2 \theta -1+2 tan \theta i}{sec^2 \theta}\\
z= \frac{\frac{sin^2 \theta}{cos^2 \theta} -1+2 \frac{sin \theta}{cos \theta}i}{\frac{1}{cos^2 \theta}}\\
z= \frac{sin^2 \theta-cos^2 \theta+2sin \theta cos \theta i}{\frac{1}{cos^2 \theta}*cos^2\theta}\\
z=- cos 2\theta + sin2 \theta i\\
Hence\\
e^{i \theta}= cos \theta + isin \theta\\
e^{-i \theta}= cos \theta - isin \theta\\
\implies e^{i2\theta}= cos 2\theta + isin 2\theta\\
\space \space \space \space \space \space \space \space \space \space e^{-i \theta}= cos 2\theta -isin2 \theta\\
Adding \space them\\
cos 2 \theta = \frac{e^{i2 \theta }-e^{-i2 \theta }}{2}\\
Also, sin 2 \theta = \frac{e^{i2 \theta }-e^{-i2 \theta }}{2i}\\
z= \frac{e^{i2 \theta }-e^{-i2 \theta }}{2}+ \frac{e^{i2 \theta }-e^{-i2 \theta }}{2i}\\
z= -e^{i2 \theta }\\
\therefore z^n= (-e^{i2 \theta })^n = (-)^ne^{i2n \theta }"